If the area of the dining room and living room is 300 sq. ft., what is the equation that would be used to solve for x? Write the equation with zero on the right, ready to solve.
Found 2 solutions by drk, richwmiller:
Answer by drk(1908)

If we assume that the dining room and living room are all in one room, then we can say
LxW = 300
which is
L x W - 300 = 0
If we cannot assume this then
L1 x W1 = dining room area
L2 x W2 = living room area
together they equal
L1 X W1 + L2 x W2 = 300
or
L1 X W1 + L2 x W2 - 300 = 0
I hope its the first option.

Answer by richwmiller(17219)

You assume too many things.
What makes you think that an equation with zero on the right is any more ready to solve than any other?
There is no equation that would be used to solve for x until you say what x will be.
So what do you want x to be?
There isn't enough info to solve for anything.